A body of mass 0.3 kg experiences gravitational force of 10 N when placed at a particular point in the field of other body of mass M. Find (a) field strength at that point and (b) acceleration due to gravity at that point.

  1. A body of mass 0.3 kg experiences gravitational force of 10 N when placed at a particular point in the field of other body of mass M. Find (a) field strength at that point and (b) acceleration due to gravity at that point.

Grade:11

1 Answers

Vikas TU
5 years ago
Dear student,
  1. Field strength is given as =  GM/r^2 (due o some other body mass M.
And Force is giveen  = 10 N
Therefore, 10 = (G*M*0.3)/r^2
Hence field strength would be => GM/r^2 where r is the distance of that point from the mass M.
 
  1. acceleration due to gravity will be = F/m => 10/0.3 = 33.33 m/s^2